OpenStudy (anonymous):

Please help simplify the expression and help explain the steps. x^2-9x+14/x^4-2x^3 It's a rational expression.

OpenStudy (anonymous):

it probably means "factor and cancel"

OpenStudy (anonymous):

Yeah, I just don't understand how to factor it out.

OpenStudy (anonymous):

even i can factor the denominator, it is \[x^4-3x^3=x^3(x-2)\]

OpenStudy (anonymous):

i suck at factoring, but my guess is the numerator has a factor of \(x-2\) as well, otherwise there is nothing to do

OpenStudy (anonymous):

since that is a good clue, it must be true that \[x^2-9x+14=(x-2)(\text{something})\]

OpenStudy (anonymous):

I get how to factor the x^2 but, it's the -9x+14 and x^4-2x3 that's confusing

OpenStudy (anonymous):

i factored \(x^4-3x^3\) for you above both terms have a common factor of \(x^3\)

OpenStudy (anonymous):

Okay, I see that.

OpenStudy (anonymous):

So how about he top?

OpenStudy (anonymous):

*the

OpenStudy (anonymous):

factor \[x^2-9x+14\] as \[(x-2)(\text{someting})\] your job is to find the "something" which should not be too hard

OpenStudy (anonymous):

okay, thanks

OpenStudy (anonymous):

but, how'd you get the 2?